hello my name is anthony and i need help with this problem cuz i am new at this and so here it is and explain it the best way you can step by step?

5(2x+3)=4(x-1)+1

Answer by rfer(16322)

5(2x+3)=4(x-1)+1
10x+15=4x-4+1
subtract 15
10x+15-15=4x-4+1-15
10x=4x-18
subtract 4x
10x-4x=4x-4x-18
6x=-18
divide by 6
6x/6=-18/6
x=-3